Data on personal consumption expenditures in the United States are shown in the table below.

Date Expenditures
May 11,674
June 11,999

Using a = 0.28 and b = 0.3, and using l0 = 11,674 and b0 = 11,999 - 11,674 as the initial values of the level and trend, respectively, develop a forecast using Holt's linear trend method for June (time period 2). Take all calculations to three decimal places. Feel free to use Excel and/or the Real Statistics app.
Answer #1

Herea = 0.28 and b = 0.3

I0 = 11674

b0 = 11999 - 11674 = 325

Here first we will evaluate trend value = b * b0  = 325 * 0.3 = 97.5

Now forecast = 11999 * 0.3 + (1 - 0.3) * 11674 = 11771.5

so

Here forecast for JUne = 11771.5 + 325 = 12096.5